# Efficient Low-Memory Fast Stack Decoding with Variance Polarization for PAC Codes

### v1

@@ VI Numerical Results
In this section, we evaluate the performance of our proposed decoding schemes over a BI-AWGN channel with BPSK modulation. Since a Fano decoder with frame error rate (FER) $D$ at a given $E_{b}/N_{0}$ achieves performance very close to the theoretical bounds for the short block lengths considered, we set the threshold parameter to $P_{\mathrm{th}}=D/10$ in our numerical results. For all evaluations, we adopt the connection polynomial $\mathbf{c}(t)=t^{10}+t^{9}+t^{7}+t^{3}+1$ [3]. Furthermore, for Fano decoding, we fix the threshold spacing to $\Delta=2$.
@@
Fig. 1 compares the performance of the conventional stack decoding algorithm for a PAC(128,64)(128,64) code with the performance of the path-pruning strategy in stack decoding (PstackD) proposed by [10] and our proposed path-pruning strategy, which leverages bit-channel variances (PstackD-Var). The average number of paths in the stack is also compared in this figure. At an $E_{b}/N_{0}$ value of 11~dB, our proposed pruning method achieves a $70\%$ reduction in the average number of paths in the stack compared to the conventional stack decoding algorithm, and an $18\%$ reduction compared to the method in [10].
TABLE I: Average Performance Metrics of PstackD-Var Decoder for PAC(64,57)(64,57) (Stack Size = 88, Max Cycles = 3232)
@@
TABLE II: Average Performance Metrics of PstackD-Var Decoder for PAC(128,99)(128,99) (Stack Size = 6464, Max Cycles = 10241024)
@@
Fig. 2 illustrates the FER performance of the PAC(64,57)(64,57) code decoded using our proposed fast stack decoding algorithm with a stack size of $8$ and a maximum of $32$ decoding cycles, compared against Fano decoding, where no limit is imposed on the number of explored paths. For comparison, we also evaluate the FER performance of the corresponding RM(64,57)(64,57) code under Viterbi, BCJR, and BP decoding as in [20]. It can be observed that while Viterbi and BCJR decoding provide near-ML performance, their computational complexity grows exponentially with the number of frozen bits, making them impractical for larger block lengths.
Table I also reports the average number of decoding cycles, the average stack size used, and the average number of times the $f$ and $g$ functions are invoked. For reference, in conventional SC decoding of polar codes, the $f$ and $g$ functions are called exactly 2N2=1262N-2=126 times for a code length of $N=64$. In contrast, for stack decoding, multiple candidate paths are maintained simultaneously, and the reported averages in the table are computed per decoded codeword, representing the total number of $f$ and $g$ function calls accumulated across all explored paths during the decoding process.
@@
Fig. 3 illustrates the FER performance of the proposed stack decoding algorithm with a stack size of $64$ and a maximum of $1024$ decoding cycles. The results demonstrate that the proposed approach achieves FER performance comparable to that of the Fano sequential decoding algorithm.

### v2

@@ VI Numerical Results
In this section, we evaluate the performance of our proposed decoding schemes over a BI-AWGN channel with BPSK modulation. To define the threshold parameter, one may use a theoretical bound, such as the RCU bound [23], denoted by $D$, and set the threshold parameter to $P_{\mathrm{th}}=D/10$. Since, for the short block lengths considered, a Fano decoder with FER $D$ at a given $E_{b}/N_{0}$ achieves performance very close to the theoretical bounds, we also can set the threshold parameter to $P_{\mathrm{th}}=D/10$ in our numerical results. The division by 10 is used to impose a more conservative performance limit. For all evaluations, we adopt the connection polynomial $\mathbf{c}(t)=t^{10}+t^{9}+t^{7}+t^{3}+1$ [15]. Furthermore, for Fano decoding, we fix the threshold spacing to $\Delta=2$.
@@
Fig. 1 compares the performance of the conventional stack decoding algorithm for a PAC(128,64)(128,64) code with an RM-based rate profile with that of the path-pruning strategy in stack decoding (PstackD) proposed in [16], which is based on the Chernoff bound, and our proposed path-pruning strategy based on bit-channel variances (PstackD-Var), which uses both the Chernoff and Chebyshev bounds. It also compares these schemes with the pruning strategy based only on the variance (Stack-Var), which relies only on the Chebyshev bound. The average number of paths in the stack (average stack used) is also shown in this figure.
At an $E_{b}/N_{0}$ value of 11~dB, our proposed pruning method achieves a $70\%$ reduction in the average number of paths in the stack compared to the conventional stack decoding algorithm, and an $18\%$ reduction compared to the method in [16].
TABLE I: Average Performance Metrics of PstackD-Var Decoder for PAC(64,57)(64,57) (Stack Size = 88, Max Cycles = $32$)
@@
TABLE II: Average Performance Metrics of PstackD-Var Decoder for PAC(128,99)(128,99) (Stack Size = $64$, Max Cycles = $1024$)
@@
Fig. 2 illustrates the FER performance of the PAC(64,57)(64,57) code with an RM-based rate profile decoded using our proposed fast stack decoding algorithm with a stack size of $8$ and a maximum of $32$ decoding cycles, compared against Fano decoding, where no limit is imposed on the number of explored paths. For comparison, we also evaluate the FER performance of the corresponding RM(64,57)(64,57) code under Viterbi, Bahl–Cocke–Jelinek–Raviv (BCJR), and belief-propagation (BP) decoding as in [1]. It can be observed that while Viterbi and BCJR decoding provide near-maximum-likelihood (ML) performance, their computational complexity grows exponentially with the number of frozen bits, making them impractical for larger block lengths.
Table I also reports the average number of decoding cycles, the average stack size used, and the average number of times the $f$ and $g$ functions are invoked. For reference, in conventional successive-cancellation (SC) decoding of polar codes, the $f$ and $g$ functions are called exactly 2N2=1262N-2=126 times for a code length of $N=64$. In contrast, for stack decoding, multiple candidate paths are maintained simultaneously, and the reported averages in the table are computed per decoded codeword, representing the total number of $f$ and $g$ function calls accumulated across all explored paths during the decoding process.
@@
Fig. 3 illustrates the FER performance of the PAC(128,99)(128,99) code with an RM-based rate profile decoded uing our proposed stack decoding algorithm with a stack size of $64$ and a maximum of $1024$ decoding cycles. The results demonstrate that the proposed approach achieves FER performance comparable to that of the Fano sequential decoding algorithm.
@@
Figure 4: FER performance comparison of PAC(64,32)(64,32) codes with Fano and stack decoding algorithms.
Fig. 4 illustrates the FER performance of a PAC(64,32)(64,32) code whose rate profile is obtained by a genetic-algorithm-based optimization initialized from the RM rate profile [14]. The code is decoded using our proposed stack decoding algorithm with a stack size of $256$ and a maximum of $1024$ decoding cycles, and its performance is compared with that of Fano decoding without any limitation on the search effort. The results demonstrate that the proposed approach achieves FER performance comparable to that of the Fano sequential decoding algorithm.
TABLE III: Average Performance Metrics of PstackD-Var Decoder for PAC(64,32)(64,32) (Stack Size = $256$, Max Cycles = $1024$)
$E_{b}/N_{0}$ (dB) 1.0 1.5 2.0 2.5 3.0 3.5 4.0 4.5 5.0 5.5
Average Cycles 63.57 51.13 33.73 25.04 20.02 17.52 16.61 16.27 16.13 16.07
Average Stack Used 78.81 60.27 44.54 33.79 26.51 23.01 20.93 20.09 19.25 18.82
Average $f$ & $g$ Operations 113.08 91.77 61.21 45.96 37.06 32.59 30.97 30.39 30.18 30.09
Table III also presents the average number of decoding cycles, the average stack size used, and the average number of times the $f$ and $g$ functions are invoked for the PAC(64,32)(64,32) code. As expected, these complexity measures decrease as $E_{b}/N_{0}$ increases. For reference, in conventional SC decoding of polar codes, the $f$ and $g$ functions are called exactly 2N2=1262N-2=126 times for a code length of $N=64$.